The recoded temperature in an engine is a r.v.X whose pdf is given by
f(x) = 
n(1 − x)
n−1
, : 0 < x < 1
0, : x < 0
where n ⩾ 1 is a known integer
(a) Show that f is indeed a pdf.
[3 Marks]
(b) Determine the corresponding cdf
1
Expert's answer
2021-09-07T02:50:58-0400

(a)


"\displaystyle\int_{-\infin}^{\infin}f(x)dx=\displaystyle\int_{0}^{1}n(1-x)^{n-1}dx"

"=[-(1-x)^n]\begin{matrix}
 1 \\
 0
\end{matrix}=-(0-1)=1, n\geq1"

Therefore "f(x)" is indeed a pdf.


(b)


"F(x)=\displaystyle\int_{-\infin}^xf(t)dt"

"0<x<1"


"F(x)=\displaystyle\int_{0}^xn(1-t)^{n-1}dt=[-(1-t)^n]\begin{matrix}
 x \\
 0
\end{matrix}"

"=1-(1-x)^n"

"F(x)=\begin{cases}
 0 &x<0\\
 1-(1-x)^n & 0\leq x<1\\
1 & x\geq 1
\end{cases}"